OpenStudy (anonymous):

how do u make a number into a fraction example 45.83

OpenStudy (anonymous):

where is that?

OpenStudy (anonymous):

Any finite decimal can be written as the whole number + decimal/ 10^number of decimal places. Your number is 45 85/100 On the other hand, if it's a recurring decimal 45.83... then you can either spot that this is 45 5/6 or you can put 45 to one side and say r = 0.83.... 10r = 8.3... Therefore 9r = 10r - r = 7.5 r = 7.5/9 = 15/18 = 5/6 Therefore 45.83... = 45 5/6 Phew - same answer as before.